OpenStudy (anonymous):

What is the equation of the line in slope intercept form that passes through the point (2,4) and is perpendicular to the line y = -2x + 4?

OpenStudy (tkhunny):

Do you have the Point-Slope Form? You'll need it.

OpenStudy (anonymous):

no really confused with this

OpenStudy (anonymous):

perpendicular line has slope that is the negative reciprocal of the slope of the given function. then substitute the point in y=mx+b

OpenStudy (anonymous):

soo... -1/2

OpenStudy (anonymous):

positive 1/2 because it is already -2

OpenStudy (anonymous):

oohh alright next???...

OpenStudy (anonymous):

substitute it (2,4) and the slope that you get in y=mx+b

OpenStudy (anonymous):

4=1/2x + ?

OpenStudy (anonymous):

solve for b